OpenStudy (anonymous):

solve (1/2)(x^2)-8=0

mathslover (mathslover):

\[\large{\frac{1}{2}*x^2=\frac{1*x^2}{2}=\frac{x^2}{2}}\] \[\large{\frac{x^2}{2}-8=0}\]

OpenStudy (lgbasallote):

so what do you get here @vmsdance ?

mathslover (mathslover):

right what r u getting from here ?

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

so i suppose you know how to proceed?

OpenStudy (anonymous):

yes